OpenStudy (anonymous):

Solve the following system of equations. Enter the y-coordinate of the solution. Round your answer to the nearest tenth. 5x+2y=21 -2x+6y=-34

Directrix (directrix):

5x+2y=21 --> Multiply each term of the first equation by (-3) -2x+6y=-34

Directrix (directrix):

-15x -6y = -63 -2x +6y =-34 Add like terms. The two y terms should add to zero. ===========

Directrix (directrix):

Then, solve for x and post what you get. Okay? @kimmykaht

OpenStudy (anonymous):

I got -5.70 @Directrix , so would the answer be -6?

OpenStudy (anonymous):

Math is really hard for me :(

Directrix (directrix):

For y, I got -64/17. Let me see what that is as a decimal to the nearest tenth.

OpenStudy (anonymous):

is it -3.8?

Directrix (directrix):

-15x -6y = -63 -2x +6y =-34 =========== -17x = - 97 x = 97/17 but the question is to find y.

Directrix (directrix):

y = - 3.8, yes, that is correct.

OpenStudy (anonymous):

thank you!! @Directrix

Directrix (directrix):

You are welcome.